OpenStudy (anonymous):

For this literal equation, find 'a'. b+c=10/a

OpenStudy (anonymous):

a(b+c)=10 Divide both sides by (b+c) a=10/(b+c)

OpenStudy (anonymous):

ok! Thx so much! :)

OpenStudy (anonymous):

Simplifying a(b + c) = 10 (b * a + c * a) = 10 (ab + ac) = 10 Solving ab + ac = 10 Solving for variable 'a'. Move all terms containing a to the left, all other terms to the right. Reorder the terms: -10 + ab + ac = 10 + -10 Combine like terms: 10 + -10 = 0 -10 + ab + ac = 0 The solution to this equation could not be determined.

OpenStudy (anonymous):

your welcome!!!!!